The length of the rectangle is 3 centimeters greater than the width. If the perimeter equals 54 centimeters, find the dimensions of the rectangle.
Answer by Maths68(1474)

The length of the rectangle is 3 centimeters greater than the width
l = w + 3 ........(1)
Perimeter = 2l + 2w
54=2l+2w
Put the value of l in above equation
54=2(w+3)+2w
54=2w+6+2w
54=4w+6
54-6=4w
48=4w
Divide by 4 both sides of above equation
48/4=4w/4
12=w
w=12
Put the value of w in (1)
l = w + 3
l = 12 + 3
l = 15
Dimensions of rectangle
Length = 15 cm
Width = 12 cm
Perimeter = 54 cm